OpenStudy (anonymous):

can someone help me out please solve for x: 2/5 + 3/5x = x+5/10

OpenStudy (jlastino):

Multiply both sides by 10 10(2/5 + 3/5x) = (x+5/10)10 4 + 6x= x+5 5x=1 x=1/5

OpenStudy (anonymous):

thanks :]